Shatin Intl. has 9.69.6 million​ shares, an equity cost of capital of 12.5 %12.5% and is expected to pay a total dividend of $ 19.2$19.2 million each year forever. It announces that it will increase its payout to shareholders. Instead of increasing its​ dividend, it will keep it constant and will start repurchasing $ 10.1$10.1 million of stock each year as well. What is your estimate of​ Shatin's stock price after this​ announcement?
Answer #1

Before announcement Price = (19.2/9.6)/.125= 16

After announcement price = (( 19.2 + 10.1)/9.6)/.125= 24.417

So due to additional cash in hand of investors, the price is expected to go up